OpenStudy (anonymous):

System B moving with Vo elastically Collides with system A at rest as Shown.At maximum Compression of Spring K.Velocity of M is?

OpenStudy (anonymous):



OpenStudy (ghazi):

conservation of momentum

OpenStudy (anonymous):

But that Did nt work..)

OpenStudy (anonymous):

Hmm, can't see the entire diagram though. Does it means that the first spring collided with the second while compressing to max?

OpenStudy (anonymous):



OpenStudy (anonymous):

@ghazi @Vincent-Lyon.Fr @JFraser @phi

OpenStudy (ghazi):

transfer whole kinetic energy of the system with mass 2m and 3m to the system with mass m and 2m , you will get it :)

OpenStudy (anonymous):

1/2 5M Vo^2 = 1/2 3M v^2

OpenStudy (anonymous):

But i think....Spring Constant has Something to Play..)

OpenStudy (ghazi):

yes spring is storing energy so you have to calculate that energy too and after that use the rule of collision, using conservation of energy and momentum. and don't forget to calculate stored potential energy of spring

OpenStudy (anonymous):

After collision assuming it is perfectly eastic the 2M will have a velocity of V. The Velocity of centre of mass is 2V/3. Now consider the motion of system A in centre of mass reference frame. When the spring is under maximum condition the velocities of both M, 2M will be zero in xcentr of mass reference frame. Hence V(absolute of M) = V(w.r.t to C.M) + V(velocity of Centr of mass) =>\[ V_{M}=V_{M,CM} + V_{CM}\] \[V_{M} = 0 + 2V/3 =2V/3\]

OpenStudy (vincent-lyon.fr):

I have never solved that kind of problems. I would start with the initial 2M-2M collision when the springs do not act yet. There would be exchange of KE and velocities of these two masses. Only then would springs start to compress.